OpenStudy (anonymous):

(2x+14)(x+14)=952 solve for x

OpenStudy (anonymous):

(2x+14)(x+14) – (2x)(x) = 952 thats really the equations

OpenStudy (anonymous):

can u help me (2x+14)(x+14) – (2x)(x) = 952

OpenStudy (pottersheep):

Expand it out

OpenStudy (pottersheep):

then bring over everything to one side making it all equal to zero

OpenStudy (pottersheep):

then you use the quadratic formula

OpenStudy (pottersheep):

When you expand it it becomes 2x^2 + 42x +101 = 0

OpenStudy (pottersheep):

Now use the quadratic formula

jimthompson5910 (jim_thompson5910):

(2x+14)(x+14) = 2x(x+14) + 14(x+14) (2x+14)(x+14) = 2x(x)+2x(14) + 14(x)+14(14) (2x+14)(x+14) = 2x^2+28x + 14x+196 (2x+14)(x+14) = 2x^2+42x+196 -------------------------------------------------- So (2x+14)(x+14) turns into 2x^2+42x+196

OpenStudy (anonymous):

x = -3/2 (7+sqrt(217))

OpenStudy (pottersheep):

I got 2x^2 + 42x +101 = 0?

jimthompson5910 (jim_thompson5910):

you're forgetting about the – (2x)(x) or -2x^2 portion

OpenStudy (anonymous):

so thats the answer jim_thompson5910 or pottersheep

OpenStudy (anonymous):

Solve for x over the real numbers: (x+14) (2 x+14) = 952 Write the quadratic polynomial on the left hand side in standard form. Expand out terms of the left hand side: 2 x^2+42 x+196 = 952 Write the quadratic equation in standard form. Divide both sides by 2: x^2+21 x+98 = 476 Solve the quadratic equation by completing the square. Subtract 98 from both sides: x^2+21 x = 378 Take one half of the coefficient of x and square it, then add it to both sides. Add 441/4 to both sides: x^2+21 x+441/4 = 1953/4 Factor the left hand side. Write the left hand side as a square: (x+21/2)^2 = 1953/4 Eliminate the exponent on the left hand side. Take the square root of both sides: x+21/2 = (3 sqrt(217))/2 or x+21/2 = -(3 sqrt(217))/2 Look at the first equation: Solve for x. Subtract 21/2 from both sides: x = (3 sqrt(217))/2-21/2 or x+21/2 = -(3 sqrt(217))/2 Look at the second equation: Solve for x. Subtract 21/2 from both sides: Answer: | | x = (3 sqrt(217))/2-21/2 or x = -21/2-(3 sqrt(217))/2

jimthompson5910 (jim_thompson5910):

so (2x+14)(x+14) – (2x)(x) = 952 turns into 2x^2+42x+196 – 2x^2 = 952 42x+196 = 952 keep going to solve for x

OpenStudy (anonymous):

so next u do 952-196

jimthompson5910 (jim_thompson5910):

good, giving you 42x = 756

OpenStudy (anonymous):

than u divide 756/42

jimthompson5910 (jim_thompson5910):

so x = ???

OpenStudy (anonymous):

18

OpenStudy (anonymous):

thank u jim

OpenStudy (anonymous):

x= 3/2 (sqrt(217)-7)

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank u jim

jimthompson5910 (jim_thompson5910):

you're welcome